OpenStudy (anonymous):

use double-angle,power reducing,half-angle to solve. TAN 112.5°

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\(112.5^\circ\) is half of \(225^\circ\) so you can use the half angle formula with \(225^\circ\)

OpenStudy (misty1212):

\[\tan(11.25^\circ)=\frac{1-\cos(225^\circ)}{\sin(225^\circ)}\] would work

OpenStudy (anonymous):

yes thank you !

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]